# Neural Flow Diffusion Models: Learnable Forward Process for Improved Diffusion Modelling

### v1

@@ Appendix C Implementation details
CIFAR-10 ImageNet 32 ImageNet 64
Channels 256 256 192
Depth 2 3 3
Channels multipliers 1,2,2,2 1,2,2,2 1,2,3,4
Heads 4 4 4
Heads Channels 64 64 64
Attention resolution 16 16,8 32,16,8
Dropout 0.0 0.0 0.0
Effective Batch size 256 1024 2048
GPUs 2 4 16
Epochs 1000 200 250
Iterations 391k 250k 157k
Learning Rate 4e-4 1e-4 1e-4
Learning Rate Scheduler Polynomial Polynomial Constant
Warmup Steps 45k 20k -
C.1 Dequantization
@@
C.2 Parameterization
We define the parameterization of the function $F_{\varphi}$ as follows in eq. 22:
$F_{\varphi}(\varepsilon,t,\mathbf{x})$ $=\mu_{\varphi}(\mathbf{x},t)+\varepsilon\sigma_{\varphi}(\mathbf{x},t).$ (48)
In addition to inferring $F_{\varphi}$, we require access to its inverse $F^{-1}_{\varphi}$ and the logarithm of the determinant of the inverse transformation $\log|J_{F}^{-1}|$ (eq. 14). The calculation of the inverse is straightforward:
$F^{-1}_{\varphi}(\mathbf{z}_{t},t,\mathbf{x})$ $=\frac{\mathbf{z}_{t}-\mu_{\varphi}(\mathbf{x},t)}{\sigma_{\varphi}(\mathbf{x},t)}.$ (49)
The logarithm of the determinant of the inverse transformation, $\log|J_{F}^{-1}|$, can also be readily determined:
$\log|J_{F}^{-1}|=-\log|J_{F}|=-\sum_{i=1}^{D}\log\left[\sigma_{\varphi}(\mathbf{x},t)\right]_{i},$ (50)
where $D$ represents the dimensionality of the data, and $\left[\sigma_{\varphi}(\mathbf{x},t)\right]_{i}$ refers to $i$-th output of the $\sigma_{\varphi}$ function.
To ensure the constraints $g_{\varphi}\geq 0$ (eq. 12) and $\bar{\sigma}_{\varphi}\geq 0$ (eq. 24) are met, we apply the softplus function to the outputs of the neural networks.

### v2

@@ Appendix F Implementation Details
CIFAR-10 ImageNet 32 ImageNet 64 AFHQ 64
Channels 256 256 192 192
Depth 2 3 3 3
Channels multipliers 1,2,2,2 1,2,2,2 1,2,3,4 1,2,2,2
Heads 4 4 4 4
Heads Channels 64 64 64 64
Attention resolution 16 16,8 32,16,8 32,16,8
Dropout 0.0 0.0 0.0 0.0
Effective Batch size 256 1024 2048 1024
GPUs 2 4 16 16
Epochs 1000 200 250 100
Iterations 391k 250k 157k 102k
Learning Rate 4e-4 1e-4 1e-4 1e-4
Learning Rate Scheduler Polynomial Polynomial Constant Constant
Warmup Steps 45k 20k - -
Our evaluation of NFDM and NFBM includes tests on synthetic data, CIFAR-10 [30], two downsampled ImageNet [11, 66] datasets, and downsampled AFHQ [8]. To maintain consistency with baselines, we employ horizontal flipping as a data augmentation technique in training models on CIFAR-10 and ImageNet [60, 58]. For density estimation of discrete data, uniform dequantization is used (see Section F.1).
We parameterize $\hat{\mathbf{x}}_{\theta}$ (eq. 11) in the reverse process using a 5-layer perceptrons with 512 neurons in each layer for synthetic data and the U-Net architecture from [12] for images. In all experiments, a 3-layer perceptrons with 64 neurons in each layer is employed to parameterize $g_{\varphi}$ (eq. 9).
For the parameterization of $F_{\varphi}$(eq. 7) in NFDM as Gaussian, we employ a neural network identical to that used for $\hat{\mathbf{x}}_{\theta}$. The only difference is that for $F_{\varphi}$, we double the output of the last layer to parameterize both $\bar{\mu}_{\varphi}$ (eq. 52) and $\bar{\sigma}_{\varphi}$(eq. 53) using the same model (see Section B.5).
For parameterizing $F_{\varphi}$ (eq. 13) in NFDM, we adopt the same approach but double the input size of the first layer to condition the forward process on both $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ (see Section C.1).
To parameterize $F_{\varphi}$ (eq. 7) in NFDM as non-Gaussian, we apply the Glow [28] architecture tailored to the corresponding dataset to transform $\varepsilon$ into $\mathbf{z}_{t}$. To incorporate conditioning, we use a neural network identical to that of $\hat{\mathbf{x}}_{\theta}$ to generate an embedding $u(\mathbf{x},t)$. We then add this conditioning to the input of each non-linear transformation in the coupling layers.
To ensure the constraints $g_{\varphi}\geq 0$ (eq. 9) and $\bar{\sigma}_{\varphi}\geq 0$ (eqs. 53 and 58) are met, we apply the softplus function to the outputs of the neural networks.
The models were trained using the Adam optimizer with the following parameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, a weight decay of $0.0$, and $\epsilon=10^{-8}$. The training process was facilitated by a polynomial decay learning rate schedule, which includes a warm-up phase for a predefined number of training steps. During this phase, the learning rate is linearly increased from $10^{-8}$ to a peak value. After reaching the peak learning rate, it is then linearly decreased to $10^{-8}$ by the final training step. The specific hyperparameters are detailed in Table 4. Training was carried out on Tesla V100 GPUs.
F.1 Dequantization
@@
F.2 NFBM with Obstacles
In Section 6.3, we demonstrate how the concept of learning NFDM with the restrictions discussed in Section 5 can be adapted to NFBM for learning generative dynamics that avoid obstacles. For this experiment, we propose training the NFBM models with an additional penalty that has higher values at points closer to obstacles. Specifically, we applied the following objective:
$\mathcal{L}_{\textrm{obs}}=\mathbb{E}_{u(t)q(\mathbf{z}_{t})}[f(\mathbf{z}_{t})],$ (72)
where $f(\mathbf{z}_{t})$ is a penalty function. In our experiments, we use the probability density of a mixture of Gaussian distributions, truncated at some low values to 00. This truncation is necessary to ensure that the model does not push trajectories infinitely away from the obstacle.
In contrast to ${\mathcal{L}_{\mathrm{crv}}}$ proposed for NFDM-OT, which penalizes the reverse process influencing the forward process as a result, $\mathcal{L}_{\textrm{obs}}$ directly penalizes the forward process, thereby influencing the properties of the reverse process.
Appendix G Additional Results
G.1 Visualisations of Forward Trajectories of NFDM-OT
Figure 3: First coordinates of forward deterministic trajectories of NFDM-OT started from points $\mathbf{x}=(-1,-1)$ and $\mathbf{x}=(1,1)$.
In this section, we provide visualizations of the forward trajectories of NFDM-OT, as illustrated in the two-dimensional data distribution experiment shown in Figure 1(b). Figure 3 depicts the deterministic trajectories of NFDM-OT. It is easy to see that the forward process has learned some highly curved trajectories. This example highlights several key aspects of NFDM-OT and NFDM more broadly.
Firstly, this example underscores the point discussed in Section D.2: while NFDM-OT penalizes the curvature of the reverse process, which is parameterized through the forward process (see Section 3.2), it does not penalize the curvature of the forward process itself.
Secondly, this example demonstrates the importance of learning nonlinear forward processes. In attempting to adapt to the straight-line reverse process, the forward process ends up being highly curved, which in turn aids the reverse process. This emphasizes the importance of the NFDM framework, which facilitates the learning of such nonlinear forward processes that may help to learn better generative dynamics.
Finally, we note that we parameterize the forward process with a Gaussian distributions featuring learnable mean and variance (see Section B.5). Even this relatively simple parameterization significantly enhances the flexibility of diffusion models compared to conventional diffusion approaches, enabling the learning of complex forward processes.
G.2 Additional Results of NFDM and NFDM-OT on Image Datasets
Table 5: Comparison of NFDM with NFDM-OT results on density estimation tasks. We present results in terms of BPD, lower is better.
Model CIFAR10 ImageNet 32
NFDM $2.49$ $3.36$
NFDM-OT $\mathbf{2.62}$ $\mathbf{3.45}$
Table 6: Comparison of NFDM with NFDM-OT results on few-step generation. We present results in terms of FID, lower is better.
CIFAR-10 ImageNet 32
Model NFE $\downarrow$ FID $\downarrow$ NFE $\downarrow$ FID $\downarrow$
$4$ $50.12$ $4$ $57.60$
NFDM $12$ $21.88$ $12$ $24.74$
$4$ $7.76$ $4$ $6.13$
NFDM-OT $12$ $5.20$ $12$ $4.11$
In this section we provide additional results, that directly compare NFDM and NFDM-OT. This results extend Tables 1 and 2. As observed, NFDM without any additional restrictions provides a better log-likelihood estimation compared to NFDM-OT. However, as expected, NFDM exhibits much worse sample quality compared to NFDM-OT in a few steps generation. We attribute this property to the higher curvature of NFDM’s generative trajectories.
G.3 Generated Samples
In this section, we present generated samples from the NFDM, NFDM-OT, and NFBM models, trained on various datasets.
NFDM: Figure 4 shows samples from the NFDM model with Gaussian $q_{\varphi}(\mathbf{z}_{t}|\mathbf{x})$ (see Section B.5). We provide samples from NFDM trained on the CIFAR-10, ImageNet 32, and ImageNet 64 datasets.
NFDM-OT. Figure 5 displays samples from NFDM trained on ImageNet 64 dataset. We generate these samples by simulating the ODE across $12$, $4$, and $2$ steps.
NFBM. Figure 6 demonstrates the generative trajectories of the NFBM model trained on the AFHQ 64 dataset. This model starts from an image of a dog at time step $t=1$ and simulate the SDE to time step $t=0$.
(a) CIFAR-10
(b) ImageNet 32
(c) ImageNet 64
Figure 4: Generated samples from NFDM trained on various datasets.
[…]
Figure 6: Generative trajectories from NFBM trained on AFHQ 64.